Oh boy, I'm stumped as to how to proceed. We're working with systems of equations right now and this problem is really starting to tax my patience.
"$24000 is invested in three different accounts, each paying 1%, 2%, and 3% interest, respectively. After 1 year, the simple interest earned from all three accounts amounts to $440. If the account paying 1% interest contains three times the amount in the account paying 2% interest, how much was invested in the three accounts?"
Specifically, I don't know how to set this information up as a system of equations. Since there are three accounts, I think this should be a 3x3 equation, but I've only been able to come up with
x+y+z=24000, and (.01)x+(.02)y+(.03)z=440


Answer by richwmiller(17219)

You are right normally we need three equations for three unknowns.
Let's see what we know and how we can use that info
"$24000 is invested in three different accounts
x+y+z=24000,
each paying 1%, 2%, and 3% interest, respectively.
After 1 year, the simple interest earned from all three accounts amounts to $440.
(.01)x+(.02)y+(.03)z=440
If the account paying 1% interest contains three times the amount in the account paying 2% interest,
You forgot this tidbit.
x=3y,
(.01)x+(.02)y+(.03)z=440,
x+y+z=24000
Sometimes you really have to squeeze info to get the three equations. Here you missed a treasure
x=12000., y=4000., z=8000.